OpenStudy (anonymous):

Solve the system of equations by graphing 9x-7y=-49 9x=7y-49

OpenStudy (anonymous):

they are the same equation

OpenStudy (anonymous):

Do I need to change it to y=mx+b

OpenStudy (anonymous):

y=7+(9/7)x

OpenStudy (anonymous):

no look at the first equation, visualize adding \(7y\) to both sides and see that you get the second equation exactly

OpenStudy (anonymous):

i dont think you wrote the problem correctly because they are the same equation yes because thats the way you know how to graph it

OpenStudy (anonymous):

It is written correctly. I am so confused.

OpenStudy (anonymous):

look at the first equation it says \(9x-7y=-49\) now you can rearrange this equation any way you like, and it will still be the same equation for example you can write it as \(7x-7y+49=0\) by adding 49 to both sides. it is still the same equation, it just looks different

OpenStudy (anonymous):

you can also write it as \[9x=4y-49\] by adding \(4y\) to both sides. nothing has changed that is why i said the first equation is the same as the second one

OpenStudy (anonymous):

ok so then i dive 9 by both?